# Status of the CORSIKA 8 particle-shower simulation code

### v1

@@ 3 Hadronic cascades
Corsika 8 employs a modular framework that integrates state-of-the-art physics models to simulate hadronic processes with precision and adaptability. It supports a versatile suite of high-energy interaction models, including EPOS-LHC (Pierog et al., (2015)), Sibyll 2.3d (Riehn et al., (2020)), QGSJet-II.04 (Ostapchenko, (2011)), and an experimental integration of Pythia 8.3 (Bierlich et al., (2022)), catering to diverse physics scenarios. Most recent efforts on the integration of the Pythia 8.3 hadronic interaction model are discussed in Gaudu, (2024) and (cite C8+P8@ECRS2024), while in parallel, ongoing work to interface the recently updated models, EPOS LHC-R (Pierog and Werner, (2023)) and QGSJet-III (Ostapchenko, (2024)) is taking place.
The model used in Corsika 8 for low-energy interactions is FLUKA (The FLUKA Collaboration et al., (2024)). The transition threshold between the high- and low-energy models is fixed at $\sim 80$ GeV as default, but can be modified. Decay processes are supported through Sibyll 2.3d and Pythia 8.
@@ Conclusion
Recent enhancements include the implementation of the low-energy interaction model FLUKA, ongoing efforts to integrate the high-energy interaction model Pythia 8.3, consideration of photo-hadronic interactions and the Landau-Pomeranchuk-Migdal (LPM) effect, refinement of the electromagnetic cascade thinning algorithm, a fully integrated Cherenkov light calculation using GPU parallelization (Albrecht et al., (2023)), and a realistic simulation of radio emission in air accompanied by a proof-of-concept CPU parallelization (Alves et al., (2023)). Extensive validation was done against Corsika 7 and other codes; what remains is understanding differences at the 5-10% level, polishing user-level and technical aspects, such as refining simulation steering, enhancing performance, and finalizing the documentation. An expert release of Corsika 8 is expected by the end of this year.

### v2

@@ 3 Hadronic cascades
Corsika 8 employs a modular framework that integrates state-of-the-art physics models to simulate hadronic processes with precision and adaptability. It supports a versatile suite of high-energy interaction models, including EPOS-LHC (Pierog et al., (2015)), Sibyll 2.3d (Riehn et al., (2020)), QGSJet-II.04 (Ostapchenko, (2011)), and an experimental integration of Pythia 8.3 (Bierlich et al., (2022)), catering to diverse physics scenarios. Most recent efforts on the integration of the Pythia 8.3 hadronic interaction model are discussed in Gaudu, (2024) and Gaudu et al., (2024), while in parallel, ongoing work to interface the recently updated models, EPOS LHC-R (Pierog and Werner, (2023)) and QGSJet-III (Ostapchenko, (2024)) is taking place.
The model used in Corsika 8 for low-energy interactions is FLUKA (Ballarini et al., (2024)). The transition threshold between the high- and low-energy models is fixed at $\sim 80$ GeV as default, but can be modified. Decay processes are supported through Sibyll 2.3d and Pythia 8.
@@ Conclusion
Recent enhancements include the implementation of the low-energy interaction model FLUKA, ongoing efforts to integrate the high-energy interaction model Pythia 8.3, consideration of photo-hadronic interactions and the Landau-Pomeranchuk-Migdal (LPM) effect, refinement of the electromagnetic cascade thinning algorithm, a fully integrated Cherenkov light calculation using GPU parallelization (Albrecht et al., (2023)), and a realistic simulation of radio emission in air accompanied by a proof-of-concept CPU parallelization (Alves et al., (2023)). Extensive validation was done against Corsika 7 and other codes; what remains is understanding differences at the 5-10% level, polishing user-level and technical aspects, such as refining simulation steering, enhancing performance, and finalizing the documentation. An expert release of Corsika 8 is expected early 2025.